HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

UPDATING THE CARBON DIOXIDE LINE LIST FOR THE HITEMP DATABASE

\underline{R. J. HARGREAVES} , I. E. GORDON, L. S. ROTHMAN, Center for Astrophysics | Harvard and Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA; X. HUANG, NASA Ames Research Center, Moffett Field, CA 94035, USA;

SETI Institute, 339 Bernado Avenue, Mountain View, CA 94043, USA

The HITEMP database^{1} (www.hitran.org/hitemp ) was established to model the radiative transfer of high temperature gaseous environments and provides accurate line-by-line molecular spectroscopic parameters for eight molecules: H_{2}O, CO_{2}, N_{2}O, CO, CH_{4}, NO, NO_{2}, and OH. The addition and update of HITEMP line lists relies on the accuracy and completeness of state-of-the-art ab initio works. To account for molecular absorption at high temperatures, it is often necessary for these works to consider billions of individual transitions per isotopologue. This makes implementing full ab initio line lists in line-by-line calculations impractical for applications such as calculating the radiative transfer of stellar, brown dwarf, and exoplanetary atmospheres. When adding methane to the HITEMP database, a methodology^{2} was developed to convert the intensity contribution of numerous weak transitions into a relatively small number of “effective” lines, thereby substantially reducing the size of the line list while retaining the absorption contribution.

The recently-published AI-3000K infrared line list for hot carbon dioxide^{3} contains over 36 billion transitions for the four most abundant CO_{2} isotopologues, and has been shown to be accurate at high temperatures. AI-3000K provides an opportunity to extend the effective line approach developed for CH_{4} in order to update the HITEMP line list of CO_{2}. This presentation will outline the development of the effective line methodology for CO_{2}, which reduces the total number of lines by 2 orders of magnitude. The CO_{2} effective lines retain some quantum information that enables line broadening parameters to be applied to the HITEMP line list. The presentation will also include detailed validations against a series of high temperature experimental spectra.

1Rothman, et al. JQSRT, 111, 2139 (2010), doi:10.1016/j.jqsrt.2010.05.001

2Hargreaves, et al. ApJS, 247, 55 (2020), doi:10.3847/1538-4365/ab7a1a

3Huang, et al. JQSRT, 392, 111748 (2023), doi:10.1016/j.jms.2023.111748 p-number: p242  Submitted on Thu, 22 Jun 2023 16:41:44 +0200